# F5R-TTS: Improving Flow Matching based Text-to-Speech with Group Relative Policy Optimization

### v1

@@ Abstract
We present F5R-TTS, a novel text-to-speech (TTS) system that integrates Gradient Reward Policy Optimization (GRPO) into a flow-matching based architecture. By reformulating the deterministic outputs of flow-matching TTS into probabilistic Gaussian distributions, our approach enables seamless integration of reinforcement learning algorithms. During pretraining, we train a probabilistically reformulated flow-matching based model which is derived from F5-TTS with an open-source dataset. In the subsequent reinforcement learning (RL) phase, we employ a GRPO-driven enhancement stage that leverages dual reward metrics: word error rate (WER) computed via automatic speech recognition and speaker similarity (SIM) assessed by verification models. Experimental results on zero-shot voice cloning demonstrate that F5R-TTS achieves significant improvements in both speech intelligibility (relatively 29.5% WER reduction) and speaker similarity (relatively 4.6% SIM score increase) compared to conventional flow-matching based TTS systems. Audio samples are available at https://frontierlabs.github.io/F5R.
Figure 1: We conduct zero-shot voice cloning experiments among three distinct models with different datasets. The assessment is performed from two key perspectives: speaker similarity (measured by SIM) and semantic accuracy (measured by WER). Higher SIM and lower WER indicate better performance.
@@ 1 Introduction
In recent years, significant advancements in Text-to-Speech (TTS) systems have enabled the generation of high-fidelity, natural-sounding voices and zero-shot voice cloning capabilities.These developments encompass both autoregressive (AR) [1, 2, 3, 4] and non-autoregressive (NAR)[5, 6, 7] model architectures. AR models typically encode audio into discrete tokens using speech codecs and then employ language model (LM)-based autoregressive models to predict these tokens, but they struggle with inference latency and exposure bias. In contrast, NAR models based on denoising diffusion or flow matching, leverage parallel computation for fast inference, demonstrating strong application potential.
Additionally, exemplified by the DeepSeek series [8, 9, 10, 11], reinforcement learning (RL) has triggered a trend in the research of large language model (LLM). RL methods like Direct Preference Optimization (DPO)[12] and Group Relative Policy Optimization (GRPO)[8] have proven effective in aligning LLM outputs with human preferences, enhancing the safety and utility of generated text through feedback optimization. In the field of image generation, RL methods like denoising diffusion policy optimization (DDPO)[13] have also been successfully applied. Currently this paradigm has extended to AR TTS systems: Seed-TTS[14] achieved RL integration using speaker similarity (SIM) and Word Error Rate (WER) as rewards with Proximal Policy Optimization (PPO)[15], REINFORCE[16] and DPO. In some other works of AR architecture, DPO and its variants are also explored[17, 18, 19, 20]. However, integrating RL into NAR architectures remain challenging due to fundamental structural divergences from LLMs. Current research shows no successful RL integration cases for NAR-based TTS, indicating that this challenge still awaits viable research solutions.
In this paper, we introduce F5R-TTS which is a novel TTS system adapting GRPO to flow-matching models through two key innovations. First, we reformulate the deterministic outputs of a flow-matching based model into probabilistic sequences, where F5-TTS[7] is adopted as the backbone for our modifications. By this way, RL algorithm application could be seamless integrated in the subsequent phase. Second, a GRPO-driven enhancement stage is designed with WER and SIM as the reward metrics, both of which are high related to human perception. Experimental validation confirms the system’s effectiveness, demonstrating measurable improvements in both intelligibility (relatively 29.5% WER reduction) and speaker consistency (relatively 4.6% SIM increase) over conventional NAR TTS baselines.
@@
  • We propose a method to transform the outputs of flow-matching based TTS models into probabilistic representations, which enables the convenient application of various reinforcement learning algorithms to flow matching models.
  • The GRPO method is successfully applied to the NAR TTS model, employing WER and SIM as reward signals.
@@ 2 Proposed Method
The proposed method divides the training process into two phases. We first pretrain the model with flow matching loss and subsequently improve the model with GRPO. In this section, we will provide a detailed explanation of how the GRPO strategy is leveraged to improve the flow matching-based model.
@@ 2.1 Preliminaries
Our model is designed mainly following F5-TTS[7], which is a novel flow matching based TTS model with zero-shot ability. The model is trained on the text-guided speech-infilling task. According to the concept of flow matching, the goal is to predict $x_{1}-x_{0}$ with $(1-t)x_{0}+tx_{1}$ as input where $x_{1}\sim$ data distribution $q(x)$ and $x_{0}\sim N(0,1)$. The vanilla objective function is defined as
@@
And we aim to further enhance the performance of the model with GRPO, which simplifies the structure based on PPO by eliminating the value model, and receives rewards by using rules or models. The penalty term KL divergence $\mathbb{D}_{KL}$ between $\pi_{\theta}$ and $\pi_{ref}$ is estimated as formula. 2.
@@ 2.2 Output Probabilization and Pretraining
We probabilize the output of the proposed model to make it better suited to GRPO and predict the distribution probability of $x_{1}-x_{0}$. Fig. 2 shows the overall structure of the model. In the first phase, we still use the flow matching objective function. The flow matching objective is to match a probability path from a standard normal distribution to a distribution approximating the data distribution.
The proposed model is trained on the infilling task as well. During training, the model takes the flow step $t$, the noisy acoustic feature $(1-t)x_{0}+tx_{1}$, the masked acoustic feature $(1-m)\odot x_{1}$, and the transcript text of complete speech $T_{gt}$ as inputs. We use the extracted mel spectrogram feature as an acoustic feature for training and pad the text feature to the same length as the acoustic feature.
@@ 2.3 Model Enhancement with GRPO
In the second phase, we further train the pretrained model as policy model $\pi_{\theta}$. At the same time, we initialize reference model $\pi_{ref}$ with the parameters of the pretrained model. The reference model is frozen throughout the whole process of GRPO. The forward operation of our TTS model during GRPO training is different from the pretraining phase. There is a sampling operation that is similar to inference. Policy model $\pi_{\theta}$ takes $x_{0}\sim N(0,1)$ as input, and then the output probability is calculated for each flow step. The sampling result of policy model $o$ is used to calculate the reward and the KL loss compared with the reference model result $x_{ref}$.
In terms of reward metrics, we select WER and SIM as the primary criteria to improve semantic consistency and speaker similarity, as these are the two most critical aspects in voice cloning tasks. We employ ASR model to transcribe the synthesized speech, obtain the transcribed text $T_{pol}$, and then calculate the WER by comparing it with the ground truth text $T_{gt}$ derived from the ground truth speech. Additionally, we utilize speaker encoder $Enc_{spk}$ to extract synthesized speaker embedding $emb_{pol}$ and ground truth speaker embedding $emb_{gt}$ from the generated speech $o$ and ground truth speech sample $x_{gt}$ respectively. The cosine similarity between these embeddings is subsequently computed to assess speaker similarity.
@@
After getting the reward, we can get the advantage through group relative advantage estimation
@@
In addition, in order to maintain the stability of the model output, GRPO also needs to use reference model $\pi_{ref}$ to provide constraints. Finally, we define formula. 10 as the objective function in the second phase.
@@ 3.1 Dataset and Experiment Setup
For the pretraining phase, we utilized WenetSpeech4TTS Basic[21], a Mandarin open-source dataset consisting of 7,226 hours of multi-speaker corpus, as the training set. During the GRPO phase, we randomly selected 100 hours of speech data from the same dataset to serve as the training set. In the evaluation, following the test setting of Seed-TTS, we synthesized 2,020 general samples and 400 hard samples using the reference speech from the Seed-TTS-eval test-cn set [^1] https://github.com/BytedanceSpeech/seed-tts-eval. General samples utilize plain text, whereas hard samples employ text that presents difficulties, such as tongue twisters or containing a high frequency of repetitive words and phrases. Additionally, we selected 70 noisy utterances from the Seed-TTS-eval test-cn set as reference utterances to generate 140 samples for testing the robustness of the model to noise.
Our model architecture was built mainly following the configuration described in the F5-TTS paper, with modifications applied only to the last output layer. During the pretraining phase, the model was trained for 1 million updates on 8 A100 40GB GPUs with a batch size of 160,000 frames. In the GRPO training phase, the model was trained for 1,100 updates on 8 A100 40GB GPUs with a batch size of 6,400. For the GRPO training, we utilized SenseVoice[22] as the automatic speech recognition (ASR) system to compute the $Reward_{W}$ and WeSpeaker[23] as the speaker encoder for $Reward_{S}$ calculation.
We selected vanilla F5-TTS as the baseline for our experiments. To demonstrate the effectiveness of GRPO in improving TTS models, we compared the performance of vanilla F5-TTS, output-probabilized F5-TTS, and the proposed method. We named them F5, F5-P, and F5-R respectively. The model architecture and parameter setting of F5 align with the original configuration. F5-P is also the pretrained model of the GRPO phase. All models were trained on the same dataset used for pretraining.
@@ 3.2.1 Visualization Analysis
We first used t-SNE [24] to visualize the speaker similarity in the 2D space. T-SNE can cluster data, such as speaker embeddings, in an unsupervised manner. In this part, we randomly chose 10 unseen speakers from Seed-TTS-eval test-cn set as target speakers, and then the models respectively synthesized 10 utterances for each speaker. WeSpeaker was utilized to get speaker embeddings of utterances. Finally, we used t-SNE to visualize the embedding. We can intuitively see the similarity between the synthesized results and the ground truth samples. We can also observe the differences between the corresponding results of different target speakers. As shown in Fig.4, the results of F5-R were well clustered according to the target speaker. Meanwhile, the subgraph of F5 and F5-P shows that the synthesis results corresponding to some target speakers were not completely clustered together. This means that the synthesis results of F5-R have better speaker similarity.
Secondly, we used the global variance (GV) [25]. GV is a method to visualize the spectral variance distribution of utterances. We respectively generated 20 converted utterances for 4 unseen speakers (2 females and 2 males) of the Seed-TTS-eval test-cn set. Then we calculated the GV of ground truth utterances and synthesized utterances. The closer the curve of the synthesized results is to the curve of ground truth utterances, the better. As shown in Fig. 5, the red curve for F5-R is more consistent with the blue curve for ground truth than others, which also shows that the synthesized results of F5-R are more similar to ground truth utterances.
@@ 3.2.2 Metrics Analysis
To evaluate the performance of the models, we computed WER and SIM as objective metrics, based on seed-tts-eval test-cn. For the calculation of these objective metrics, we utilized the testing tools provided by seed-tts-eval. Specifically, for WER, we employed Paraformer-zh[26] to transcribe. For SIM, we utilized a WavLM-large-based speaker verification model[27] to extract speaker embedding. WER and SIM separately reflect the models’ performance in terms of semantic accuracy and speaker similarity. WER is desired to be as low as possible, whereas SIM is preferred to be as high as possible. Table. 1 shows the result of objective evaluation for the three models on the testset.
In terms of SIM, it can be observed that F5 and F5-P exhibit comparable performance across two distinct testsets, with F5 marginally outperforming the latter. Our model, however, demonstrates superior performance on both testsets, emerging as the top-performing model. Notably, on the general testset, our model surpasses other models by a minimum of 0.03 SIM score. Compared to F5, F5-R achieved a 4.6% and 5.6% relative increase on the general testset and the hard testset respectively. This indicates that GRPO contributes positively to enhancing speaker similarity.
In terms of WER, F5 and F5-P remain closely aligned across the two distinct testsets. Our model, however, demonstrates superior performance on both testsets. On the general testset, our model achieves a WER relative reduction of 29.5% compared to the baseline. The advantage of the proposed model is also pronounced on the hard testset, where it exhibits a WER relative reduction of 5.9%. This further underscores the efficacy of GRPO in enhancing semantic accuracy.
For the noise robustness test, we used noisy utterances as reference audio. We found that noise had little effect on the SIM performance of the three models, but caused an increase in the WER of F5 and F5-P. F5-R showed better robustness under noise conditions. Meanwhile, we found that F5-R also had better rhythm performance under noise conditions. Audio samples of noise robustness test are available at https://frontierlabs.github.io/F5R.
@@
To demonstrate the generalizability of the proposed method, we conducted identical experiments on additional datasets. For the pretraining phase, we employed an internal Mandarin dataset comprising 10,000 hours of corpus, primarily sourced from radio broadcasts and audiobooks. We conducted a preliminary quality-based filtering for the dataset. Furthermore, we randomly selected a 100-hour subset from this collection to serve as the trainset for the GRPO phase. The objective metrics comparison with the internal dataset is shown in Table. 2. Compared to F5, F5-R achieved a relatively 18.4% WER reduction and a relatively 3.1% SIM increase on the general testset. On the hard testset, F5-R achieved a WER relative reduction of 8.1% and a SIM relative increase of 1.1%. On the noisy testset, F5-R achieved a WER relative reduction of 26.1% and a SIM relative increase of 2.2%. The overall results align with those obtained on WenetSpeech4TTS Basic, showing measurable improvements of all models across in terms of objective metrics. Our experiments demonstrate that GRPO consistently enhances model performance across diverse datasets.
Overall, the performance of F5 and F5-P is largely comparable. As anticipated, the GRPO with WER and SIM as rewards enables the model to achieve gains in both semantic accuracy and speaker similarity. Guided by the speaker-related component of the reward, the model is better able to clone the characteristics of the target speaker through in-context learning. On the hard testset, the relative advantage of the proposed model in terms of WER is even more pronounced. We hypothesize that this may be attributed to the WER-related component of the reward, which enhances the model’s ability to preserve semantic content more effectively. However, the performance of all three models deteriorates on the hard testset, indicating that increased text complexity generally leads to a decline in model stability. This observation may serve as a focal point for future optimization efforts.

### v2

# F5R-TTS: Improving Flow-Matching based Text-to-Speech with Group Relative Policy Optimization

@@ Abstract
We present F5R-TTS, a novel text-to-speech (TTS) system that integrates Gradient Reward Policy Optimization (GRPO) into a flow-matching based architecture. By reformulating the deterministic outputs of flow-matching TTS into probabilistic Gaussian distributions, our approach enables seamless integration of reinforcement learning algorithms. During pretraining, we train a probabilistically reformulated flow-matching based model which is derived from F5-TTS with an open-source dataset. In the subsequent reinforcement learning (RL) phase, we employ a GRPO-driven enhancement stage that leverages dual reward metrics: word error rate (WER) computed via automatic speech recognition and speaker similarity (SIM) assessed by verification models. Experimental results on zero-shot voice cloning demonstrate that F5R-TTS achieves significant improvements in both speech intelligibility (a 29.5% relative reduction in WER) and speaker similarity (a 4.6% relative increase in SIM score) compared to conventional flow-matching based TTS systems. Audio samples are available at https://frontierlabs.github.io/F5R.
Figure 1: We conducted zero-shot voice cloning experiments comparing three distinct models across different datasets. The evaluation was performed from two key perspectives: speaker similarity (measured by SIM) and semantic accuracy (measured by WER). Higher SIM and lower WER indicate superior performance.
@@ 1 Introduction
In recent years, significant advancements in Text-to-Speech (TTS) systems have enabled the generation of high-fidelity, natural-sounding voices and zero-shot voice cloning capabilities.These developments span both autoregressive (AR)  [1, 2, 3, 4] and non-autoregressive (NAR) [5, 6, 7] model architectures. AR models typically encode audio into discrete tokens using speech codecs and then employ language model (LM) based autoregressive models to predict these tokens. However, such approaches suffer from inference latency and exposure bias. In contrast, NAR models based on denoising diffusion or flow matching, leverage parallel computation for faster inference speeds, demonstrating strong application potential.
Additionally, as exemplified by the DeepSeek series  [8, 9, 10, 11], reinforcement learning (RL) has triggered a trend in the research of large language model (LLM). RL methods like Direct Preference Optimization (DPO) [12] and Group Relative Policy Optimization (GRPO) [8] have proven effective in aligning LLM outputs with human preferences, enhancing the safety and utility of generated text through feedback optimization. In the field of image generation, RL methods like denoising diffusion policy optimization (DDPO) [13] have also been successfully applied. This paradigm has now extended to AR TTS systems: Seed-TTS [14] achieved RL integration using speaker similarity (SIM) and Word Error Rate (WER) as rewards with Proximal Policy Optimization (PPO) [15], REINFORCE [16] and DPO. In some other works of AR architecture, DPO and its variants are also explored [17, 18, 19, 20]. However, integrating RL into the NAR architectures remains challenging due to fundamental structural divergences from LLMs. Current research shows no successful cases of RL integration in NAR based TTS systems, indicating that this challenge still awaits viable research solutions.
In this paper, we introduce F5R-TTS which is a novel TTS system adapting GRPO to flow-matching models through two key innovations. First, we reformulate the deterministic outputs of a flow-matching based model into probabilistic sequences, where F5-TTS [7] is adopted as the backbone for our modifications. This reformulation enables seamless integration of RL algorithms in the subsequent phase. Second, a GRPO-driven enhancement stage is designed with WER and SIM as the reward metrics, both of which are highly correlated with human perception. Experimental results demonstrate the system’s effectiveness, showing significant improvements in both speech intelligibility (a 29.5% relative reduction in WER) and speaker consistency (a 4.6% relative increase in SIM score) compared to conventional NAR TTS baselines.
@@
  • We propose a method to transform the outputs of flow-matching based TTS models into probabilistic representations, which enables convenient application of various reinforcement learning algorithms to flow-matching models.
  • We successfully apply the GRPO method to NAR-TTS models, using WER and SIM as reward signals.
@@ 2 Proposed Method
The proposed method divides the training process into two phases. We first pretrain the model with flow matching loss and subsequently improve the model with GRPO. In this section, we will provide a detailed explanation of how the GRPO strategy is leveraged to improve the flow-matching based model.
@@ 2.1 Preliminaries
Our model is designed mainly following F5-TTS [7], which is a novel flow-matching based TTS model with zero-shot ability. The model is trained on the text-guided speech-infilling task. According to the concept of flow matching, the goal is to predict $x_{1}-x_{0}$ with $(1-t)x_{0}+tx_{1}$ as input where $x_{1}\sim$ data distribution $q(x)$ and $x_{0}\sim N(0,1)$. The vanilla objective function is defined as
@@
And we aim to further enhance the performance of the model using GRPO, a simplified variant of PPO that eliminates the value model and computes rewards via rule-based or model-based methods. The penalty term KL divergence $\mathbb{D}_{KL}$ between $\pi_{\theta}$ and $\pi_{ref}$ is estimated as shown in formula. 2.
@@ 2.2 Output Probabilization and Pretraining
We reformulate the model’s output into probabilistic terms to enhance compatibility with GRPO, enabling prediction of the distribution probability for $x_{1}-x_{0}$. Fig. 2 shows the overall structure of the model. In the first phase, we retain the flow matching objective function. The flow matching objective is to match a probability path from a standard normal distribution to a distribution approximating the data distribution.
The proposed model is trained on the infilling task as well. During training, the model takes the flow step $t$, the noisy acoustic feature $(1-t)x_{0}+tx_{1}$, the masked acoustic feature $(1-m)\odot x_{1}$, and the transcript text of complete speech $T_{gt}$ as inputs. We use extracted mel-spectrogram feature as acoustic feature for training and pad the text feature to the same length as the acoustic feature.
@@ 2.3 Model Enhancement with GRPO
In the second phase, we further train the pretrained model as policy model $\pi_{\theta}$ while initializing the reference model $\pi_{ref}$ with the pretrained parameters. The reference model remains frozen throughout GRPO phase. The forward operation of our TTS model during GRPO training is different from the pretraining phase. There is a sampling operation that is similar to inference. Policy model $\pi_{\theta}$ takes $x_{0}\sim N(0,1)$ as input, and then the output probability is calculated for each flow step. The sampling result of policy model $o$ is used to calculate rewards and the KL loss compared with the reference model result $x_{ref}$.
In terms of reward metrics, we select WER and SIM as the primary criteria to improve semantic consistency and speaker similarity, as these represent the two most critical aspects in voice cloning tasks. We employ an ASR model to transcribe the synthesized speech, obtain the transcribed text $T_{pol}$, and then compare the transcriptions with the ground truth text $T_{gt}$ of the ground truth speech to calculate WER. Additionally, we utilize a speaker encoder $Enc_{spk}$ to extract synthesized speaker embedding $emb_{pol}$ and ground truth speaker embedding $emb_{gt}$ from the generated speech $o$ and ground truth speech sample $x_{gt}$ respectively. Speaker similarity is evaluated by computing cosine similarity between these embeddings.
@@
After calculating the reward, we can get the advantage through group relative advantage estimation [8].
@@
In order to maintain the stability of the model output, GRPO also needs to use the reference model $\pi_{ref}$ to provide constraints. Finally, we define formula. 10 as the objective function in the second phase.
@@ 3.1 Dataset and Experiment Setup
For the pretraining phase, we utilized WenetSpeech4TTS Basic [21], a Mandarin open-source dataset consisting of 7,226 hours of multi-speaker corpus, as the training set. During the GRPO phase, we randomly selected 100 hours of speech data from the same dataset for training. In the evaluation, following the test setting of Seed-TTS, we synthesized 2,020 general samples and 400 hard samples using the reference speech from the Seed-TTS-eval test-cn set [^1] https://github.com/BytedanceSpeech/seed-tts-eval. General samples utilize plain text, whereas hard samples employ text that presents difficulties, such as tongue twisters or containing a high frequency of repetitive words and phrases. To test noise robustness, we generated 140 samples using 70 noisy utterances from the same test set.
Our model architecture was built mainly following the configuration described in the F5-TTS paper, with modifications applied only to the last output layer. During the pretraining phase, the model was trained for 1 million updates on 8 A100 40GB GPUs with a batch size of 160,000 frames. In the GRPO training phase, the model was trained for 1,100 updates on 8 A100 40GB GPUs with a batch size of 6,400 frames. For the GRPO training, we utilized SenseVoice [22] as the automatic speech recognition (ASR) system to compute the $Reward_{W}$ and WeSpeaker [23] as the speaker encoder for $Reward_{S}$ calculation.
We selected vanilla F5-TTS as the baseline for our experiments. To demonstrate the effectiveness of GRPO in improving TTS models, we compared the performance of vanilla F5-TTS, output-probabilized F5-TTS, and the proposed method. We named them F5, F5-P, and F5-R respectively. The F5 strictly preserves the original architecture and parameter settings. F5-P is also the pretrained model of the GRPO phase. We trained all models on the identical pretraining dataset.
@@ 3.2.1 Visualization Analysis
We first used t-SNE [24] to visualize the speaker similarity in the 2D space. T-SNE can cluster data, such as speaker embeddings, in an unsupervised manner. For this analysis, we randomly selected 10 unseen speakers from Seed-TTS-eval test-cn set as target speakers, and then the models respectively synthesized 10 utterances per speaker. WeSpeaker was utilized to get speaker embeddings of utterances, which were then visualized via t-SNE. We can intuitively see the similarity between the synthesized results and the ground truth samples. We can also observe distribution differences across target speakers. As shown in Fig.4, the results of F5-R were well clustered according to the target speaker. Meanwhile, the subgraph of F5 and F5-P shows that the synthesis results corresponding to some target speakers were not completely clustered together. This means that the synthesis results of F5-R have better speaker similarity.
Secondly, we used the global variance (GV) [25]. GV is a method to visualize the spectral variance distribution of utterances. We respectively generated 20 converted utterances for 4 unseen speakers (2 females and 2 males) of the Seed-TTS-eval test-cn set. We then calculated GV for both reference and synthesized utterances. A closer match between the synthesized and reference curves indicates better performance. As shown in Fig. 5, the red curve for F5-R shows better alignment with the blue curve for reference utterances than others, which also shows that the synthesized results of F5-R are more similar to reference utterances.
@@ 3.2.2 Metrics Analysis
To evaluate model performance, we adopted WER and SIM as objective metrics, based on seed-tts-eval test-cn. For metric computation, we leveraged the official evaluation toolkit provided with seed-tts-eval. For WER, we employed Paraformer-zh [26] to transcribe. For SIM, we utilized a WavLM-large-based speaker verification model [27] to extract speaker embedding. These metrics respectively quantify semantic accuracy (lower WER preferred) and speaker similarity (higher SIM desired). Table. 1 presents the comparative evaluation results across the three models.
In terms of SIM, it can be observed that F5 and F5-P exhibit comparable performance across two distinct test sets, with F5 marginally outperforming the latter. Our proposed model achieves superior performance on both sets, establishing itself as the top performer. Notably, on the general test set, our model outperforms others with a minimum advantage of 0.03 SIM points. Compared to F5, F5-R achieved a 4.6% and 5.6% relative increase on the general test set and the hard test set respectively. This indicates that GRPO contributes positively to improving speaker similarity.
In terms of WER, F5 and F5-P remain closely aligned across the two distinct test sets. Our model, however, achieves significantly better results on both sets. On the general test set, it attains a 29.5% relative WER reduction compared to the baseline, with a further 5.9% reduction observed on the hard test set. These results conclusively confirm the effectiveness of GRPO in enhancing semantic accuracy.
For the noise robustness test, we used noisy utterances as reference audio. Noise exerted negligible influence on SIM across all models, while inducing an increase in the WER of F5 and F5-P. Meanwhile, F5-R demonstrated significantly improved noise robustness. And we found that F5-R also maintained superior rhythm performance under noise conditions. Audio samples of noise robustness test are available at https://frontierlabs.github.io/F5R.
@@
To demonstrate the generalizability of the proposed method, we conducted parallel experiments on additional datasets. For the pretraining phase, we employed an internal Mandarin dataset comprising 10,000 hours of corpus, primarily sourced from radio broadcasts and audiobooks. We conducted a preliminary quality-based filtering for the dataset. Furthermore, a 100-hour subset was randomly selected for GRPO training. The objective metrics comparison with the internal dataset is shown in Table. 2. Compared to F5, F5-R achieved an 18.4% relative reduction in and a 3.1% relative increase in SIM on the general test set. On the hard test set, F5-R achieved a WER relative reduction of 8.1% and a SIM relative increase of 1.1%. On the noisy test set, F5-R achieved a WER relative reduction of 26.1% and a SIM relative increase of 2.2%. The overall results align with those obtained on WenetSpeech4TTS Basic, showing that GRPO consistently enhances model performance across diverse datasets.
Overall, the performance of F5 and F5-P remains largely comparable. As anticipated, the GRPO with WER and SIM as rewards enables the model to achieve gains in both semantic accuracy and speaker similarity. Guided by the speaker-related component of the reward, the model demonstrates enhanced capability to clone target speaker characteristics through in-context learning. On the hard test set, the proposed model exhibits more pronounced relative advantages in WER performance. We hypothesize that this improvement stems from the WER-related reward component, which effectively strengthens the model’s semantic preservation ability. However, all three models show performance degradation on the hard test set, suggesting that increased text complexity generally reduces model stability. This observation may serve as a key focus for future optimization efforts.